Reports 1-1 of 1 Clear search Modify search
AdV-SGD (FDS commissioning)
delauren, m'ishid, ding, - 22:32 Wednesday 22 June 2022 (56219) Print this report
Green and IR Alignment Recovery after the EQB! sorbotane replacement

Tuesday afternoon activities

After the EQB1 sorbotane dumper replacement (56211), we checked the bench leveling and being in the our specificaton, we start with the green alignment recovery.

The strategy was the following

  • To find the retroreflection from the FC, since just before the sorbotane replacement the cavity was well aligned, we did not touched the mirrors on the benches after EQB1, supposing that are well aligned;
  • we centered SQB2_M1 and M2 by acting on EQB1_M5 and M7. then we centered FCIM  with SQB2_M1 and we observed has we expected some reflection from the FCIM:
  • At that point we aceted on SQB1_M1 to center the beam on FCIM and to overlapp it on the incoming beam: we had the retroreflection on EQB1_GR_cam and GR_PD.

We where not able to check the beam on FCEM because there was a filter in front to the FCEB_GR_cam.

Wednesday Morning Activitiy: FC locked on the green

We start by removing from FCEB both the diagphram and the filter respectively in front to FCEB_GR_cam and GR_PD to acces to the full window of both (see picture attached).

  •   we centered better FCIM by looking to the FCIM front camera ( http://ipcam53.virgo.infn.it),;
  •  then we cacted on FCIM marionett to  align the reflected beam on the incoming beam;
  • once we have again the beam on EQB1_GR_PD and GR_cam; we misaligned FCIM  (-1000 mrad far from the previous position) to look at the its trasmition and center the FCEM looking on the camera that looks it from the cavity side  (cavity side) ttp://ipcam58.virgo.infn.it in order to have the transmitted beam on FCEB_GR_cam;
  • we move the FCEM marionettes to overlapp the beam reflected to the incoming one looking to SQB1_M1 and M2 from the tank windows;
  • at the end we replace FCIM and we find the resonance and we start to align the cavity.

At this point we had some problem with FC logitudinal  looking  loops even all the signals looked fine. At the end we asked to Mateusz and Marco from remote to give a look and Marco  solved the problem  by following the procedure in the comment.

Note that @ 10.00 UTC  we had a problem with FCIM_LC (drfit of all the DoF, (see  Menzione report 56224).We asked to Alessandro  to give a look and we suggest us to resatrt the FCIM_LC process and close again all the loop on marionette (after the loop restart we open and close each of the three loops)  and we solved the problem.

Wednesday Afternoon Activity:IR alignment recovery

We followed a procedure similar to the one followed for the green

  • We start from the recovery on SQB1: with RR on, we look for the SC reflection  from it to EQB1 by mooving EQB1_5 and 7 by hand (we unload all the M3 and M5 actuator by setting them to 0 V);
  • once we had the beam on HD_M6, M5 and M4 we worked to theoverlap it to the LO on HD_cams first by continuing to act on EQB1_M5 and M7 and then on HD_M6 and M4.
  • being sure in this way that the beam was passing through the Double Faraday in SQB1, we switched off the RR and looked to the EQB2_IR_cam and PD: as aspected we have a siignal on them and worked to maximize them
  • once we had a good signal  (spot without cut on EQB2) we start to look for the IR resonance, looking the FCEB_IR_cam and PD and by acting on SQB1_M21 and M23 and we find them.

REMAING WORK:

-> Align the SC on OPO (it is misaligned and we have to check if this is consequence of the OPO misalignemnt due to the humidity wilde fluctuation or to our action on the bench, even we worke just on the M5-M6/ HD zone - and we opened just thes sector).

-> Maximization of the SC resonance;

-> check the GR power (at the present the maximum transmission id ~8.9V )

At this point since  we note that the SC was misaligned we decide to stop and continue tomorrow morning, since in any case the planned activities on stray ligh loop cannot be done due to the problem with the OPO lockiing.

 

Images attached to this report
Search Help
×

Warning

×